The Sanchez family is determined to purchase a $305,000 home without incurring any debt. The family plans to save $5,600 a quarter for this purpose and expects to earn 6.65 percent, compounded quarterly. How long will it be until the family can purchase a home?
Answer #1

Pq nor 6 Answert Given datar • The Sanchez family is determined to purchase a $ 305,000 home. . -The family plans to save $→ [((1+(0.0665/4)?)] = 0.905 4687499, → 1.0166259 = 0.9054687499 +i -1.9054687499 n = log 1-9054687499 / Jog 1.016625 = 0.280